Walt made an extra $9000 last year from a part-time job. He invested part of the money at a 9% annual rate of interest, and the rest at an 8% annual rate of interest. At the end of the year, he made a total of $770 in interest. How much was invested at 8%?
Answer by josmiceli(19441)

Let x = the amount invested at 9%
Since the total amount invested was $9000, the amount invested at 8%
can be expressed as (9000 - x)
the interest received from the 9% investment is
.09x
the interest received from the 8% investment is
.08(9000 - x)
If you add these two interest incomes together, you get the toal interest
that Walt made, so
.09x + .08(9000 - x) = 770
.09x + 720 - .08x = 770
.01x = 50
x = 5000
this is the amount invested at 9%
(9000 - 5000)= 4000 is the amount invested at 8%